Solve for and give the voltage across the 2 Ohm resistor in the frequency domain. W 12 -j4 12/ 0°u 4 /90° A
Answer #1

2 vo 1220V 0 42904 roll 2-2 = ? Detron, Salutron Apply nodol of point v V/220 + ju = 4270=4290 > V-12 + V (2 tju! 2-14)(2+0 45 -12 + VC2+jų) = 4490 20 > SV-60 + 2v + Vijų - 4496 20 3 067+34) = 80290+60 V = 12:40 12